### v1

@@ 3. $a=3$
where $0\leqslant\alpha_{u},\alpha_{v},\beta_{u},\beta_{v}\leqslant x$; we denote them by $u_{long}^{+\alpha,\beta}$ and $v_{long}^{+\alpha,\beta}$ correspondingly. So, $0\leqslant a-b\leqslant 2$. Let $u_{short}^{-\alpha,\beta}$ be the subword of $u$ which contain $n_{1}$ occurrences of 1 and all small blocks (all blocks except $0^{x}$, $0^{\alpha}$ and $0^{\beta}$). In other words, $u_{short}^{-\alpha,\beta}$ is obtained from $u_{short}$ by removing two blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$. The word $v_{short}^{-\alpha,\beta}$ is defined symmetrically.
@@
Consider the subword $v^{\prime}=1^{n_{1}}0^{x}$ of $v$. We now overlay by 1’s the word $v$ on $u$ in a way that $v^{\prime}$ coincides with an equal subword in $u$. Since $u$ and $v$ are not equal, for each such overlay we can find and add to $v^{\prime}$ block of 0’s in $v$ which is bigger than the corresponding block in $u$. Adding such blocks of 0’s to $v^{\prime}$, we get a subword $v^{\prime\prime}$ of $v$ which is not a subword of $u$, and $|v^{\prime\prime}|\leqslant n_{1}+(a+1)x$. So either lemma is proved, or $n_{1}+(a+1)x\geqslant\frac{3}{4}n+5$. So, in the rest of the proof we will assume that
$n_{1}+(a+1)x\geqslant\frac{3}{4}n+5\quad\text{and}\quad n_{1}+(b+1)x\geqslant\frac{3}{4}n+5$ (6)
We say that $u_{short}^{-\alpha,\beta}$ has a period with $d<n_{1}$ occurrences of 1’s if there exist positive integers $p_{1},\dots,p_{i}$ and $q_{1},\dots,q_{i}$ such that $u_{short}^{-\alpha,\beta}=(1^{p_{1}}0^{q_{1}}\ldots 1^{p_{i}}0^{q_{i}})^{\frac{n_{1}}{d}}$, where $p_{1}+\ldots+p_{i}=d$.
We can now describe the plan of the rest of the proof:
@@
Case 2. $u_{short}^{-\alpha,\beta}\ncong v_{short}^{-\alpha,\beta}$ or $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$ and $u_{short}^{-\alpha,\beta}$ is either non-periodic or periodic but there is no period with $ks$ occurrences of 1’s.
2.1. $l\geqslant\frac{n}{4}-(b-2)(x-2)$.
2.2. $l<\frac{n}{4}-(b-2)(x-2)$.
Case 3. $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$ and $u_{short}^{-\alpha,\beta}$ is periodic with a period containing $ks$ occurrences of 1’s.
3.1. $s\geqslant 2$.
3.2. $s=1$.
3.2.1. $\alpha_{u}>\alpha_{v}$.
3.2.2. $\alpha_{u}=\alpha_{v},\beta_{u}=\beta_{v}$
Case 2. $u_{short}^{-\alpha,\beta}\ncong v_{short}^{-\alpha,\beta}$ or $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$ and $u_{short}^{-\alpha,\beta}$ is either non-periodic or periodic but there is no period with $ks$ occurrences of 1’s.
@@ 3.2.2. $\alpha_{u}=\alpha_{v},\beta_{u}=\beta_{v}$
Without loss of generality we assume that $|u_{short}^{-\alpha,\beta}|\leqslant|v_{short}^{-\alpha,\beta}|$ (the proof in the case $|u_{short}^{-\alpha,\beta}|\geqslant|v_{short}^{-\alpha,\beta}|$ is symmetric).
2.1. $l\geqslant\frac{n}{4}-(b-1)(x-2)+2$.
@@
Consider the subword $v_{1}=1^{n_{1}}0^{x}$ of $v$ and fix some its occurrence in $v$. There are at most $a$ occurrences of the subword $v_{1}$ in $u$. For each of these $a$ occurrences we can overlay by 1’s the word $v$ on $u$ such that the chosen occurrence of $v_{1}$ in $v$ coincides with the occurrence of $v_{1}$ in $u$. Recall that we can assume that $dist(u_{long}^{+\alpha,\beta})=dist(v_{long}^{+\alpha,\beta})$, since the case of inequality has been considered earlier (Case 1).
@@
If $u_{short}^{-\alpha,\beta}\ncong v_{short}^{-\alpha,\beta}$, then for each overlay there is a block of 0’s in $v$ which is, first, not one of big blocks $0^{x}$ or one of the two blocks $0^{\alpha}$ and $0^{\beta}$ we removed from $u_{short}$, and, secondly, which is bigger than the corresponding block in $u$. We now add all such blocks to $v_{1}$, and we let $v_{1}^{\prime}$ denote the obtained subword of $v$.
@@
If $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$ and $u_{short}^{-\alpha,\beta}$ does not have a period with $ks$ occurrences of 1’s, then for one overlay the subwords $u_{short}^{-\alpha,\beta}$ and $v_{short}^{-\alpha,\beta}$ coincide (if they coincide for more than one overlay, then they have a period with $ks$ occurrences of 1’s). For each overlay for which they do not coincide, we add blocks of 00’s to $v_{1}$ in the same way as in the case $u_{short}^{-\alpha,\beta}\ncong v_{short}^{-\alpha,\beta}$. For the overlay for which $u_{short}^{-\alpha,\beta}$ and $v_{short}^{-\alpha,\beta}$ coincide, one of the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ is smaller than the corresponding block in $v$. We then add block $0^{\alpha_{u}+1}$ or $0^{\beta_{u}+1}$ to $v_{1}$ to the corresponding place. In this case we let $v_{1}^{\prime}$ denote the obtained subword of $v$.
@@
The word $v_{1}^{\prime}$ is the subword of the word $v$ and is not a subword of $u$. When we constructed $v_{1}^{\prime}$, we did not add at least $l-b-a$ small blocks. So, we added to $v_{1}$ (of length $n_{1}+x$) at most $n_{short,v}-(l-a-b)+1$ zeros. Indeed, if $u_{short}^{-\alpha,\beta}\ncong v_{short}^{-\alpha,\beta}$, then we added zeros to $v_{1}$ from only small blocks. If $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$, then $a-1$ times we added zeros from small blocks which is not equal to $0^{\alpha_{v}}$ and $0^{\beta_{v}}$, and one time we added the block $0^{\alpha_{u}+1}$ or $0^{\beta_{u}+1}$ which is less then $10^{\alpha_{v}+1}+$ 1 or $0^{\beta_{v}+1}$ since $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$. So,
$|v_{1}^{\prime}|\leqslant(n_{1}+x)+n_{short,v}-(l-a-b-1)\leqslant$
since $2\geqslant a-b\geqslant 0$, then
$\leqslant n_{1}+n_{short,v}+bx-bx-l+2b+x+3\leqslant$
now since $n_{1}+n_{short,v}+bx=|v|=n$, we have
$\leqslant n-l-(b-1)(x-2)+5\leqslant\frac{3}{4}n+3,$
@@
where the latter inequality comes from the condition $l\geqslant\frac{n}{4}-(b-1)(x-2)+2$ of the case 2.1.
2.2. $l<\frac{n}{4}-(b-1)(x-2)+2$.
@@
Consider the subword $u_{zeros}$ of $u$, which contains $n_{1}$ occurrences of 1 and one 0 from each block of 0’s, and fix some its occurrence in $u$. Then either $v$ does not have a subword which is equal to $u_{zeros}$ or there is a subword $v_{zeros}=u_{zeros}$ of $v$. The first case is simple: the word $u_{zeros}$ is the subword of $u$ and not a subword of $v$ and $|u_{zeros}|\leqslant\frac{3}{4}n+4$ (it will follow from further inequalities). In the second case we proceed depending on the form of the word $u_{zeros}$. Fix some occurrence $v_{zeros}$ in $v$.
@@
If $u_{zeros}$ is not periodic, then there is only one way to take $u_{zeros}$ in $u$ modulo the selection of 0’s from each block of 0’s. We now overlay by 1’s the word $v$ on $u$ such that such that $v_{zeros}$ and $u_{zeros}$ coincide. Since $v\neq u$, then there is a block $0^{s}$ in $v$ which is bigger than the corresponding block of 0’s in $u$. We add this block to $v_{zeros}$ to the corresponding place. Let $v^{\prime}$ denote the obtained subword of $v$. The word $v^{\prime}$ is the subword of the word $v$ and is not a subword of $u$, and its length at most
$|v^{\prime}|\leqslant|v|+(x-1)=n_{1}+l+(x-1)\leqslant$
now we use the inequality $l<\frac{n}{4}-(b-1)(x-2)+2$ from the condition of the case 2.2.:
$\leqslant n_{1}+\frac{n}{4}-(b-1)(x-2)+x+1=n_{1}+\frac{n}{4}-(b-2)(x-2)+3\leqslant\frac{3}{4}n+3,$
@@
where the last inequality follows from the inequalities $b\geqslant 2$, $x\geqslant 2$ and $n_{1}\leqslant\frac{n}{2}$. In particular, we have $|u_{zeros}|\leqslant\frac{3}{4}n+4$.
@@
If $u_{zeros}$ is periodic, then $u_{zeros}=(1^{\alpha_{1}}01^{\alpha_{2}}0\ldots 1^{\alpha_{s}}0)^{\frac{l}{s}}$. Consider the word $u_{1}=(1^{\alpha_{1}}01^{\alpha_{2}}0\ldots 1^{\alpha_{s}}0)(10)^{l-s}$ and fix some its occurrence in $u$. The word $u_{1}$ has at most $\frac{n+l}{2}$ occurrences of 1 and $l$ occurrences of 0 (proved in Lemma 4.2, case 2.1.2). Then either $v$ does not have a subword which equal to $u_{1}$ or there is a subword $v_{1}=u_{1}$ of $v$. In the first case $u_{1}$ is the subword of the word $u$ and not a subword of $v$, and we will show that $|u_{1}|\leqslant\frac{3}{4}n+4$. In the second case we add one block $0^{x}$ to $v_{1}$. We let $v_{1}^{\prime}$ denote the obtained subword of $v$ and we fix its occurrences in $v$. Then there are at most $a$ overlays by 1’s of $v$ on $u$ such that $v_{1}^{\prime}$ coincides with an equal subword. For each overlay we can add blocks as in the Case 2.1 depending on whether $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$ or not. We let $v_{1}^{\prime\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime\prime}$ has at most $\frac{n_{1}+l}{2}$ occurrences of $1$, one block $0^{x}$, one 00 from each of the remaining $b-1$ blocks $0^{x}$ and at most $n_{short}$ added 00’s from other blocks of 0’s. Now, by inequality (6), we have
$n_{short,v}=n_{0}-bx=(n_{0}+n_{1}+x)-(n_{1}+(b+1)x)\leqslant n+x-\frac{3}{4}n-5=\frac{n}{4}-5+x.$ (7)
So, $v_{1}^{\prime\prime}$ is a subword of $v$ and not a subword of $u$ and its length is at most
$\frac{n_{1}+l}{2}+x+n_{short,v}+b-1\leqslant\frac{n_{1}+n_{short,v}+bx}{2}-\frac{bx}{2}+\frac{l}{2}+\frac{n_{short,v}}{2}+x+b-1\leqslant$
@@
Since $\frac{n_{1}+n_{short,v}+bx}{2}=n$, by (7) and due to inequality $l<\frac{n}{4}-(b-1)(x-2)+2$ from the condition of the case 2.2. we have
@@
$=\frac{3}{4}n-\frac{2(b-2)(x-2)}{2}+\frac{1}{2}\leqslant\frac{3}{4}n+\frac{1}{2},$
since $b\geqslant 2$ and $x\geqslant 2$. In particular, $|u_{zeros}|\leqslant\frac{3}{4}n+4.$
@@
Case 3. $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$ and $u_{short}^{-\alpha,\beta}$ is periodic with the period containing $ks$ occurrences of 1’s. In other words, there exist positive integers $p_{1},\dots,p_{i}$ and $q_{1},\dots,q_{i}$ such that $u_{short}^{-\alpha,\beta}=(1^{p_{1}}0^{q_{1}}\ldots 1^{p_{i}}0^{q_{i}})^{(\frac{n_{1}}{ks})}$, where $p_{1}+\ldots+p_{i}=ks$.
@@
Since $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$, $dist(u)=dist(v)$ and $n_{0,u}=n_{0,v}$, then $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$. So, if one of the words $u_{long}$ and $v_{long}$ is a word of the first type, then the other word is a word of the first type as well. Then both words have period with $ks$ occurrences of 1’s, and since $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$, we have $u=v$, which gives a contradiction. Hence $u_{long}$ and $v_{long}$ are words of the second or the third type. In particular, $0\leqslant a-b\leqslant 1$. Consider two cases depending on $s$.
@@
3.1. $s\geqslant 2$. We consider two subcases depending on the types of the words $u_{long}$ and $v_{long}$. The proofs of the two cases are similar; however, we provide details for both.
@@
(a) At least one of the words $u_{long}$ and $v_{long}$ is a word of the second type.
@@
Assume that $u_{long}$ is a word of the second type, that is, in $u_{long}^{+\alpha,\beta}$ we have $\alpha<k,\beta=k$. Then $n_{1}=k(a+1)$. Consider the subword $u_{1}=(1^{p_{1}}0^{q_{1}}\cdots 1^{p_{i}}0^{q_{i}})1^{n_{1}-ks}$ of $u$. There are at most $\frac{k(a+1)}{ks}\leqslant\frac{a+1}{2}$ subwords of $u_{short}^{-\alpha,\beta}$ which are equal to $u_{1}$. Then either $v_{short}^{-\alpha,\beta}$ does not have a subword which is equal to $u_{1}$ or there is a subword $v_{1}=u_{1}$ of $v$. In the first case we add one block $0^{x}$ to $u_{1}$. Let $u_{1}^{\prime}$ denote the obtained subword of $u$. The word $u_{1}^{\prime}$ is a subword of the word $u$ and is not a subword of $v$, and we will later prove that $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+4$.

### v2

@@ 4.2 Lemmas constituting the proof
where $0\leqslant\alpha_{u},\alpha_{v},\beta_{u},\beta_{v}\leqslant x$; we denote them by $u_{\operatorname{long}}^{+\alpha,\beta}$ and $v_{\operatorname{long}}^{+\alpha,\beta}$, respectively. So, $0\leqslant a-b\leqslant 2$. Let $u_{\operatorname{short}}^{-\alpha,\beta}$ be the subword of $u$ which contain $n_{1}$ occurrences of 1 and all small blocks (all blocks except $0^{x}$, $0^{\alpha}$ and $0^{\beta}$). In other words, $u_{\operatorname{short}}^{-\alpha,\beta}$ is obtained from $u_{\operatorname{short}}$ by removing two blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$. The word $v_{\operatorname{short}}^{-\alpha,\beta}$ is defined analogously.
@@
Consider the subword $v^{\prime}=1^{n_{1}}0^{x}$ of $v$. We now consider a 1-overlay of $v$ on $u$ such that $v^{\prime}$ coincides with an occurrence of an equal subword in $u$. Since $u$ and $v$ are not equal, for each such 1-overlay we can find and add to $v^{\prime}$ a block of 0’s in $v$ which is bigger than the corresponding block in $u$. Adding such blocks of 0’s to $v^{\prime}$, we get a subword $v^{\prime\prime}$ of $v$ which is not a subword of $u$, and $|v^{\prime\prime}|\leqslant n_{1}+(a+1)x$. So, either we have a desired distinguishing subword, or $n_{1}+(a+1)x\geqslant\frac{3}{4}n+5$. So, in the rest of the proof we assume that
$n_{1}+(a+1)x\geqslant\frac{3}{4}n+5\quad\text{and}\quad n_{1}+(b+1)x\geqslant\frac{3}{4}n+5.$ (6)
We can now describe the plan of the rest of the proof. We distinguish between two main cases: Case 3 treats the situation when $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$ (here we use notation from Definition 2.7) and $u_{\operatorname{short}}^{-\alpha,\beta}$ is periodic with a period containing $ks$ occurrences of 1 for some integer $s$; Case 2 treats the opposite case. Case 2 is divided into two subcases corresponding to whether inequality $l\geqslant\frac{n}{4}-(b-2)(x-2)$ holds or not. Case 3 is divided into subcases 3.1 corresponding to $s\geqslant 2$, and 3.2 corresponding to $s=1$. The latter case is further split into two subcases depending on whether $\alpha_{u}=\alpha_{v}$ or not (note that in the case of equality we also have $\beta_{u}=\beta_{v}$ due to the conditions of Case 3).
@@
Case 2. $u_{\operatorname{short}}^{-\alpha,\beta}\ncong v_{\operatorname{short}}^{-\alpha,\beta}$ or $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$ and $u_{\operatorname{short}}^{-\alpha,\beta}$ is either non-periodic or periodic, but there is no period with $ks$ blocks of 1’s for some integer $s$.
@@
Without loss of generality we assume that $|u_{\operatorname{short}}^{-\alpha,\beta}|\leqslant|v_{\operatorname{short}}^{-\alpha,\beta}|$ (the proof in the case $|u_{\operatorname{short}}^{-\alpha,\beta}|\geqslant|v_{\operatorname{short}}^{-\alpha,\beta}|$ is symmetric).
Case 2.1. $l\geqslant\frac{n}{4}-(b-1)(x-2)+2$.
@@
Consider the subword $v_{1}=1^{n_{1}}0^{x}$ of $v$ and fix some of its occurrence in $v$. There are at most $a$ ways to take an occurrence of a subword of $u$ which is equal to $v_{1}$. For each of these $a$ occurrences we can consider a 1-overlay of $v$ on $u$ such that the chosen occurrence of $v_{1}$ in $v$ coincides with the chosen occurrence of $v_{1}$ in $u$. Recall that we can assume that $dist(u_{\operatorname{long}}^{+\alpha,\beta})=dist(v_{\operatorname{long}}^{+\alpha,\beta})$, since the case of inequality has been considered earlier (Case 1).
@@
If $u_{\operatorname{short}}^{-\alpha,\beta}\ncong v_{\operatorname{short}}^{-\alpha,\beta}$, then for each 1-overlay there is a block of 0’s in $v$ which is not one of big blocks $0^{x}$ or one of the two blocks $0^{\alpha}$ and $0^{\beta}$ we removed from $u_{\operatorname{short}}$, and which is bigger than the corresponding block in $u$. We now add all such blocks to $v_{1}$, and we let $v_{1}^{\prime}$ denote the obtained subword of $v$.
@@
If $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$ and $u_{\operatorname{short}}^{-\alpha,\beta}$ does not have a period with $ks$ blocks of 1’s, then the subwords $u_{\operatorname{short}}^{-\alpha,\beta}$ and $v_{\operatorname{short}}^{-\alpha,\beta}$ coincide only for one 1-overlay (if they coincide for more than one overlay, then they have a period with $ks$ blocks of 1’s). For each 1-overlay for which they do not coincide, we add blocks of 0’s to $v_{1}$ in the same way as in the case $u_{\operatorname{short}}^{-\alpha,\beta}\ncong v_{\operatorname{short}}^{-\alpha,\beta}$. Consider a 1-overlay for which $u_{\operatorname{short}}^{-\alpha,\beta}$ and $v_{\operatorname{short}}^{-\alpha,\beta}$ coincide. For this 1-overlay one of the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ is smaller than the corresponding block in $v$. We then add a block $0^{\alpha_{u}+1}$ or $0^{\beta_{u}+1}$ to $v_{1}$ to the corresponding place, and we let $v_{1}^{\prime}$ denote the obtained subword of $v$.
@@
The word $v_{1}^{\prime}$ is the subword of the word $v$ and is not a subword of $u$. When we constructed $v_{1}^{\prime}$, we did not add at least lbal-b-a small blocks. So, we added to $v_{1}$ (of length $n_{1}+x$) at most $n_{\operatorname{short},v}-(l-a-b)+1$ zeros. Indeed, if $u_{\operatorname{short}}^{-\alpha,\beta}\ncong v_{\operatorname{short}}^{-\alpha,\beta}$, then we added zeros to $v_{1}$ from only small blocks. If $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$, then a1a-1 times we added zeros from small blocks which are not equal to $0^{\alpha_{v}}$ and $0^{\beta_{v}}$, and one time we added the block $0^{\alpha_{u}+1}$ or $0^{\beta_{u}+1}$ which is less then $0^{\alpha_{v}+1}$ or $0^{\beta_{v}+1}$ since $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$. So,
$|v_{1}^{\prime}|\leqslant(n_{1}+x)+n_{\operatorname{short},v}-(l-a-b-1).$
Since $2\geqslant a-b\geqslant 0$, we have
$(n_{1}+x)+n_{\operatorname{short},v}-(l-a-b-1)\leqslant n_{1}+n_{\operatorname{short},v}-l+2b+x+3.$
Now since $n_{1}+n_{\operatorname{short},v}+bx=|v|=n$, we have
$n_{1}+n_{\operatorname{short},v}+bx-bx-l+2b+x+3\leqslant n-l-(b-1)(x-2)+5\leqslant\frac{3}{4}n+3,$
@@
where the latter inequality comes from the condition $l\geqslant\frac{n}{4}-(b-1)(x-2)+2$ of Case 2.1. Combining this series of inequalities, we obtain $|v_{1}^{\prime}|\leqslant\frac{3}{4}n+3.$
Case 2.2. $l<\frac{n}{4}-(b-1)(x-2)+2$.
@@
Consider the subword $u_{\operatorname{zeros}}$ of $u$, which contains $n_{1}$ occurrences of 1 and one 0 from each block of 0’s, and fix some its occurrence in $u$. Then either $v$ does not have a subword which is equal to $u_{\operatorname{zeros}}$ or there is a subword $v_{\operatorname{zeros}}=u_{\operatorname{zeros}}$ of $v$. The first case is simple: the word $u_{\operatorname{zeros}}$ is a subword of $u$ and not a subword of $v$ and $|u_{\operatorname{zeros}}|\leqslant\frac{3}{4}n+4$ (we will prove this inequality later, together with the second case). In the second case we proceed depending on the form of the word $u_{\operatorname{zeros}}$. Fix some occurrence $v_{\operatorname{zeros}}$ in $v$.
@@
If $u_{\operatorname{zeros}}$ is not periodic, then there is only one way to take $u_{\operatorname{zeros}}$ in $u$ modulo the selection of 0 from each block of 0’s. We now consider the unique 1-overlay of $v$ on $u$ such that such that $v_{\operatorname{zeros}}$ and $u_{\operatorname{zeros}}$ coincide. Since $v\neq u$, there is a block $0^{s}$ in $v$ which is bigger than the corresponding block of 0’s in $u$. We add this block to $v_{\operatorname{zeros}}$ to the corresponding place. Let $v^{\prime}$ denote the obtained subword of $v$. The word $v^{\prime}$ is a subword of the word $v$ and is not a subword of $u$. We now estimate its length:
$|v^{\prime}|\leqslant|v_{zeros}|+(x-1)=n_{1}+l+(x-1)$
Now using the inequality $l<\frac{n}{4}-(b-1)(x-2)+2$ from the condition of Case 2.2, we get
$n_{1}+l+(x-1)\leqslant n_{1}+\frac{n}{4}-(b-1)(x-2)+x+1=n_{1}+\frac{n}{4}-(b-2)(x-2)+3\leqslant\frac{3}{4}n+3,$
@@
where the last inequality follows from the inequalities $b\geqslant 2$, $x\geqslant 2$ and $n_{1}\leqslant\frac{n}{2}$. In particular, we have $|u_{\operatorname{zeros}}|\leqslant\frac{3}{4}n+4$.
@@
If $u_{\operatorname{zeros}}$ is periodic, then $u_{\operatorname{zeros}}=(1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{s}}0)^{\frac{l}{s}}$ for some integer $s$. Consider the word $u_{1}=(1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{s}}0)(10)^{l-s}$ and fix an occurrence of it in $u$. The word $u_{1}$ has at most $\frac{n+l}{2}$ occurrences of 1 and $l$ occurrences of 0 (proved in Lemma 4.2, Case 2.1.2). Then either $v$ does not have a subword which is equal to $u_{1}$ or there is a subword $v_{1}=u_{1}$ of $v$. In the first case $u_{1}$ is a subword of the word $u$ and not a subword of $v$, and we show that $|u_{1}|\leqslant\frac{3}{4}n+4$. In the second case we add one block $0^{x}$ to $v_{1}$. We let $v_{1}^{\prime}$ denote the obtained subword of $v$ and fix an occurrence of it in $v$. Then there are at most $a$ 1-overlays of $v$ on $u$ such that $v_{1}^{\prime}$ coincides with an occurrence of an equal subword. For each 1-overlay we can add blocks as in Case 2.1 depending on whether $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$ or not. We let $v_{1}^{\prime\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime\prime}$ has at most $\frac{n_{1}+l}{2}$ occurrences of $1$, one block $0^{x}$, one 0 from each of the remaining b1b-1 blocks $0^{x}$ and at most $n_{\operatorname{short}}$ added 0’s from other blocks of 0’s. Now, by inequality (6), we have
$n_{\operatorname{short},v}=n_{0}-bx=(n_{0}+n_{1}+x)-(n_{1}+(b+1)x)\leqslant n+x-\frac{3}{4}n-5=\frac{n}{4}-5+x.$ (7)
@@
So, $v_{1}^{\prime\prime}$ is a subword of $v$ and is not a subword of $u$; we now estimate its length. Since $\frac{n_{1}+n_{\operatorname{short},v}+bx}{2}=\frac{n}{2}$, by (7) and due to inequality $l<\frac{n}{4}-(b-1)(x-2)+2$ from the condition of Case 2.2, we have the following:
$|v_{1}^{\prime\prime}|\leq\frac{n_{1}+l}{2}+x+n_{\operatorname{short},v}+b-1\leqslant\frac{n_{1}+n_{\operatorname{short},v}+bx}{2}-\frac{bx}{2}+\frac{l}{2}+\frac{n_{\operatorname{short},v}}{2}+x+b-1\leqslant$
@@
$\frac{3}{4}n-\frac{2(b-2)(x-2)}{2}+\frac{1}{2}\leqslant\frac{3}{4}n+\frac{1}{2},$
since $b\geqslant 2$ and $x\geqslant 2$. In particular, $|u_{\operatorname{zeros}}|\leqslant\frac{3}{4}n+4.$
@@
Case 3. $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$ and $u_{\operatorname{short}}^{-\alpha,\beta}$ is periodic with period containing $ks$ blocks of 1’s. In other words, there exist positive integers $p_{1},\dots,p_{i}$ and $q_{1},\dots,q_{i}$ such that $u_{\operatorname{short}}^{-\alpha,\beta}=(1^{p_{1}}0^{q_{1}}\cdots 1^{p_{i}}0^{q_{i}})^{\frac{n_{1}}{ks}}$, where $p_{1}+\ldots+p_{i}=ks$, for some integer $s$.
@@
Since $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$, $dist(u)=dist(v)$ and $n_{0,u}=n_{0,v}$, we have $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$. So, if one of the words $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ is a word of the first type, then the other word is a word of the first type as well. Then both words have period with $ks$ blocks of 1’s, and since $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$, we have $u=v$, which gives a contradiction. Hence $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ are words of the second or the third type. In particular, $0\leqslant a-b\leqslant 1$. Consider two cases depending on $s$.
@@
Case 3.1. $s\geqslant 2$. We consider two subcases depending on the types of the words $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$. The proofs of the two cases are similar; however, we provide details for both.
@@
Case 3.1.1. At least one of the words $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ is a word of the second type.
@@
Assume that $u_{\operatorname{long}}$ is a word of the second type, that is, in $u_{\operatorname{long}}^{+\alpha,\beta}$ we have $\alpha<k,\beta=k$. Then $n_{1}=k(a+1)$. Consider the subword $u_{1}=(1^{p_{1}}0^{q_{1}}\cdots 1^{p_{i}}0^{q_{i}})1^{n_{1}-ks}$ of $u$. There are at most $\frac{k(a+1)}{ks}\leqslant\frac{a+1}{2}$ ways to choose an occurrence of a subword of $u_{\operatorname{short}}^{-\alpha,\beta}$ which is equal to $u_{1}$. Then either $v_{\operatorname{short}}^{-\alpha,\beta}$ does not have a subword which is equal to $u_{1}$ or there is a subword $v_{1}=u_{1}$ of $v$. In the first case we add one block $0^{x}$ to $u_{1}$. Let $u_{1}^{\prime}$ denote the obtained subword of $u$. The word $u_{1}^{\prime}$ is a subword of the word $u$ and is not a subword of $v$, and we later prove that $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+4$.